# Review and Prospect of Algebraic Research in Equivalent Framework between Statistical Mechanics and Machine Learning Theory

### v2

@@ 3.1 Mathematical Results
Assume that $\varphi(w)$ is a $C_{0}^{\infty}$ class function and that $W\subset\mathbb{R}^{d}$ is a compact subset whose open kernel $W^{o}$ contains the support of $\varphi(w)$, $\mbox{supp}\;\varphi$. Let us define a function of $W$
@@
Note that $L(w)\geq S$, where $S$ is the entropy of $q(x)$. We assume that $L(w)$ is a real analytic function of $w\in W^{o}$ and that there exists $w_{0}\in\mbox{supp}\;\varphi$ which minimizes $L(w)$ $(w\in W)$. The set of all parameters which minimize $L(w)$ in $W$ is denoted by
@@
In other words, $W_{0}$ consists of all zero points of the real analytic function $L(w)-L(w_{0})$, which is called an analytic set. If $w_{0}$ satisfies $q(x)=p(x|w_{0})$, then $q(x)$ is said to be realizable by $p(x|w)$. Note that, if $q(x)$ is realizable by $p(x|w)$, then $L(w_{0})$ is equal to the entropy $S$ of $q(x)$. If there exists a unique $w_{0}$ that minimizes $L(w)$ and if the Hessian matrix $\nabla^{2}L(w_{0})$ is positive definite, then $q(x)$ is said to be regular for $p(x|w)$, if otherwise it is said to be singular for $p(x|w)$.
@@
where $w^{*}$ is the unique parameter that minimizes $H(w,X^{n})$. The conventional theory for regular case is called regular learning theory. Statistical theory for singular case is called singular learning theory which contains regular theory as a special case. If a learning machine contains latent variables or hierarchical structure, then the regularity condition is not satisfied [30, 29, 22], hence we need singular learning theory both in statistics and machine learning. For example, in deep learning, the rank of the Hessian matrix in the neighborhood of the obtained parameter by training is far smaller the dimension of $w$.
@@
We need a method how to analyze the set
@@
where $k=(k_{1},k_{2},...,k_{d})$ and $h=(h_{1},h_{2},...,h_{d})$ are multi-indices of nonnegative integers, in which at least one $k_{i}$ is a positive integer. Here $b(u)>0$ is a positive analytic function, and $|g^{\prime}(u)|$ is the absolute value of the Jacobian determinant of $w=g(u)$. The correspondence between $W\setminus W_{0}$ and $g^{-1}(W\setminus W_{0})$ is one-to-one. Note that a function $w=g(u)$ is called proper if the inverse image of a compact set is also compact. The real log canonical threshold (RLCT) $\lambda$ and multiplicity $m$ are defined by
@@
where $\min_{L.C.}$ and $\max_{L.C.}$ show the minimum and maximum values over all local coordinates respectively. Here we define $(h_{j}+1)/(2k_{j})=\infty$ for $k_{j}=0$, and ##\## means the number of elements of a set. Hence $0<\lambda<\infty$ and $1\leq m\leq d$. Since $W$ is compact and $w=g(u)$ is proper, the number of all local coordinates can be taken to be finite. The property and application of RLCT is discussed in subsection 3.2.
@@
where $f(x,w)=\log(p(x|w_{0})/p(x|w))$. Since $W$ is a compact set, both sides of eq.(13) are trivially finite. Hence eq.(13) requires the inequality for the case $f(x,w)\rightarrow 0$, in other words, the variance of $f(x,w)$ is bounded by its average. In the case when this condition is not satisfied, the variance of Hamiltonian cannot be bounded by its average, resulting that the free energy has a different asymptotic behavior. [50]. By using resolution theorem and the relatively finite condition, it is proved that the following limit is finite
@@
holds without explicit values of $\lambda$ and $\nu(\beta)$, which shows that the difference between the generalization error and the training error is asymptotically equal to the posterior variance of point-wise log likelihood function. Hence we can compare their theoretical values with the experimental ones.
@@ 3.2 Real log canonical threshold
which was firstly created by Gel’fand [23] and has been studied by many mathematicians. By the definition, $\zeta(z)$ is an holomorphic function in $\Re(z)>0$ which can be analytically continued to the unique meromorphic function on the entire complex plane. This unique analytic continuation can be proved by either the Hironaka resolution theorem in subsection 3.1 or the existence of Bernstein-Sato polynomial [14]. There exist both a differential operator $D_{w}$ and a polynomial $b(z)$ such that, for arbitrary $z\in\mathbb{C}$ and $w\in W$,
@@
Then the monic polynomial that has the lowest order and satisfies this equation is called Bernsterin-Sato polynomial. By Hironaka resolution theorem, All poles of the zeta function is real and negative numbers. The largest pole and its order of the zeta function are equal to $(-\lambda)$ and $m$ respectively. By using the zeta function, it is trivial that statistical learning theory of a pair
@@
    If there exists the unique $w_{0}\in W_{0}$ such that $\nabla^{2}L(w_{0})$ is positive definite, and $\varphi(w_{0})>0$, then $\lambda=d/2$.
@@
By using these properties, RLCTs of important statistical models and learning machines were found by developing resolution procedures in neural networks [69, 7] deep linear networks [6], deep convolution ReLU network [51] with and without skip connection, normal mixtures [85], Poisson mixtures [57], multinomial mixtures [83], general and nonnegative matrix facorizations [5, 31], Boltzmann machines [87], hidden and general Markov models [86, 92] and latent Dirichlet allocations [32].
[…]
@@
where $\hat{w}$ is the parameter which minimizes $H(w,X^{n})$. In regular cases, the difference between the free energy and BIC is a constant order random variable. For general cases when $q(x)$ may be singular for for $p(x|w)$, the singular BIC was proposed [19] by using the estimated RLCT $\hat{\lambda}$,
@@ 3.3 Singular Fluctuation
In this subsection, mathematical properties and applications of singular fluctuation are studied. Singular fluctuation appears in many statistical contexts [1, 2, 3, 48].
@@
If $q(x)$ is regular for $p(x|w)$, the limit $\beta\rightarrow\infty$ results in the maximum likelihood method in statistics. The posterior distribution converges to the delta function on the maximum likelihood estimator. Then
@@
holds, which was firstly found by Akaike [1]. The $2n$ times of the right hand side of this equation is called Akaike information criterion. If $q(x)$ is not realizable by $p(x|w)$, then $d$ is replaced by $\mbox{tr}(IJ^{-1})$ which is called by Takeuchi information criterion. For the case when $\beta=1$ and $q(x)$ is regular for and realizable by a model, the $\mathbb{E}_{X^{n}}[G(1,X^{n})]$ can be estimated by deviance information criterion (DIC) [61]. The right hand side of eq.(18) is called WAIC which can be used to estimate the generalization error in general cases.
@@
Note that $C_{n}(\beta)$ can be numerically calculated by using the posterior average,
@@
The difference between LOOCV and WAIC is smaller than any $1/n$ order random variable. In numerical calculation of eq.(22), the stability of the expectation $\mathbb{E}_{w}^{\beta}[p(X_{i}|w)^{-\beta}]$ is not ensured if a leverage sample point is contained in a sample [55, 21]. A new calculation method of LOOCV was proposed by using the approximation in MCMC distribution [64]. When the conditional probability is estimated in regression problems, LOOCV requires that the sample is independent, whereas information criteria do only conditional independence. On the other hand, information criteria needs that the sample size $n$ is sufficiently large, hence LOOCV and information criteria are different statistical tools [77, 80]. The concept of generalization error was extended onto the problems such as covariate shift, causal inference [37], and overparametrized cases [53], and the generalized information criteria are now being developed.
@@ 3.4 Singular Learning Theory and AI Alignment
In 1966, a statistician I. J. Good, who firstly found the importance of marginal likelihood $Z(1,X^{n})$ in statistical inference, speculated as [28]
@@
Artificial neural network research started in the 1960s. Nowadays, artificial intelligence (AI) research is progressing so rapidly that it is difficult to predict to which direction AI takes us. It is said that alignment of artificial intelligence is necessary for safety and welfare of human being in the near future, where AI alignment is sometimes defined as AI working without deviating from the designer’s intention.
In order to consider AI alignment, we need all opinions from all different viewpoints, and problems to be studied are beging proposed [4, 12]. Note that singularities in learning machines make the free energy and the generalization loss smaller if Bayesian inferences, hence degenerate property of hierarchical learning machines can be considered to be an advantage [84]. In AI alignment, we need to study developmental landscape of in-context learning in also transfomers [35]. Deneracy in the loss function [43] is useful in mechanistic interpretability [18].
In this subsection, we discuss AI alignment from the viewpoint of equivalence of statistical mechanics and machine learning. There are at least three problems caused by singularities in artificial neural networks: lack of identifiability, difficulty in design of the prior distribution, and phase transition in learning process. These three points originated from the hierarchical structure of learning machines. Remark that hierarchical structure is necessary for a learning machine to have an ability of universal approximation, that is to say, an arbitrary continuous function can be approximated by multi-layered neural networks with higher precision. Moreover it makes the generalization error far smaller if a learning machine is over-parametrized. Therefore, the following three problems should be solved.
The first problem is lack of identifiability. A statistical model or a leaning machine $p(x|w)$ is called identifiable if the map $w\mapsto p(x|w)$ is one-to-one. Many of classical statistical models are identifiable, hence the role of the parameter is uniquely determined. Therefore, by checking the obtained parameter, we can examine the reason why a machine answers an output for a given inputs. However, in singular learning machines, the role of the parameter is not uniquely determined, hence even by checking the parameter, we cannot examine whether the model’s learning is subject to the designer’s intention.
@@
and preparing the equivalence class $W/\sim$, the map from a parameter to probability density function can be made one-to-one. This method is called blow-down in algebraic geometry. However, if $w_{0}$ is a singularity in the original parameter space, then its neighborhood in $W/\sim$ has infinite dimension, resulting that difficulty in understanding the parameter cannot be solved by $W/\sim$.
The second problem is difficulty in design of prior distribution. In older Bayesianism in the 20th century, it was premised that uncertainty could be captured by a statistical model and the personal belief should be represented by a prior distribution. However, because of nonidentiability, each element of the parameter does not have any concrete meaning, hence the prior distribution cannot be determined by any a priori personal knowledge about data-generating process. One might think that Jeffreys’ prior may be chosen as an objective one, however, it makes the generalization error far larger than others, which destroys the advantage of hierarchical structure.
We need modern Bayesian statistics, in which a statistical model or a learning machine is only a candidate and that a prior distribution is not any belief of a person but a part of a statistical model. In the practical applications, ridge or lasso type prior distribution with hyperparameter is sometimes employed and the hyperparameter is optimized by several criteria including cross validation.
The third problem is the phase transition phenomenon in learning process. In classical statistical models, the optimal parameter $w_{0}$ is unique and the posterior distribution can be approximated as a normal distribution which converges to the delta function $\delta(w-w_{0})$ as the sample size increases. The posterior distribution may be understood as a gradual increase of confidence.
On the other hand, in deep learning, even if the sample size is huge, it is far smaller than the infinity for the deep neural network. Let $q(x)$ be unrealizable by $p(x|w)$. By using division of unity, there is a set of nonngegative $C_{0}^{\infty}$ class functions $\{\varphi_{k}(w)\}$ such that
@@
$F_{nk}(\beta)==nL(w_{k})+\frac{\lambda_{k}}{\beta}\log n+O(\log\log n)$
and
@@
Therefore the free energy is given by the minimum of the local free energy. In statistical learning theory, $L(w_{k})$ and $\lambda_{k}\log n$ are called bias and variance, respectively. Note that $\exp(-F_{nk})$ is in proportion to the posterior probability of the local coordinate. Therefore the optimal local parameter set for minimum free energy is automatically chosen by the posterior probability. However, it is not equal to the optimal parameter set for the minimum generalization error. The local parameter set chosen by the posterior distribution jumps from a neighborhood to another one as sample size invreases. Such jumping process makes difficult to diagnose whether the learning process is out of order or not.
Phase transition phenomena can be also observed in the different situation. In mixture models, a Dirichlet distribution is often chosen for a prior distribution of the mixture ratio. Then the posterior distribution has a phase transition according to the hyperparameter of Dirichlet distribution [77, 83]. In the different phases, RCLTs are different and the supports of asymptotic posterior distributions are different [77].

### v3

@@ 3.1 Mathematical Results
Assume that a prior distribution, $\varphi(w)$, is a $C_{0}^{\infty}$ class function of $\mathbb{R}^{d}$ and that $W\subset\mathbb{R}^{d}$ is a compact subset whose open kernel contains the support of $\varphi(w)$, $\mbox{supp}\;\varphi$. Let us define a function of $W$
@@
Note that $L(w)\geq S$, where $S$ is the entropy of $q(x)$. We assume that $L(w)$ is a real analytic function on the open kernel of $W$ and that there exists $w_{0}\in\mbox{supp}\;\varphi$ which minimizes $L(w)$ $(w\in W)$. The set of all parameters which minimize $L(w)$ in $W$ is denoted by
@@
In other words, $W_{0}$ consists of all zero points of the real analytic function $L(w)-L(w_{0})$, which is called an analytic set. If $w_{0}$ satisfies $q(x)=p(x|w_{0})$, then $q(x)$ is said to be realizable by $p(x|w)$. Note that, if $q(x)$ is realizable by $p(x|w)$, then $L(w_{0})$ is equal to the entropy $S$ of $q(x)$. If there exists a unique $w_{0}$ that minimizes $L(w)$ and if the Hessian matrix $\nabla^{2}L(w_{0})$ is positive definite, then $q(x)$ is said to be regular for $p(x|w)$; otherwise it is said to be singular for $p(x|w)$.
@@
where $w^{*}$ is the unique parameter that minimizes $H(w,X^{n})$ and $C$ is a normalizing constant. The conventional theory for regular cases is called regular learning theory.
Statistical learning theory that for singular case is called singular learning theory. From the mathematical point of view, singular learning theory contains regular learning theory as a very special example, hence the former is a generalization of the latter. If a learning machine contains latent variables or hierarchical structure, then the regularity condition is not satisfied [28, 29, 30], hence we need singular learning theory both in statistics and machine learning. For example, in deep learning, the rank of the Hessian matrix in the neighborhood of the obtained parameter by training is far smaller than the dimension of $w$ in general.
@@
We need a method to analyze the set
@@
where $k=(k_{1},k_{2},...,k_{d})$ and $h=(h_{1},h_{2},...,h_{d})$ are multi-indices of nonnegative integers, in which at least one $k_{i}$ is a positive integer. Here $b(u)>0$ is a positive analytic function, and $|g^{\prime}(u)|$ is the absolute value of the Jacobian determinant of $w=g(u)$. The correspondence between $W\setminus W_{0}$ and $g^{-1}(W\setminus W_{0})$ is one-to-one. Note that a function $w=g(u)$ is called proper if the inverse image of a compact set is also compact. Concerning eqs. (10) and (11), we introduce the real log canonical threshold (RLCT) $\lambda$ and multiplicity $m$ by
@@
where $\min_{L.C.}$ and $\max_{L.C.}$ denote the minimum and maximum values over all local coordinates, respectively. Here we define $(h_{j}+1)/(2k_{j})=\infty$ for $k_{j}=0$, and ##\## is the number of elements of a set. Hence $0<\lambda<\infty$ and $1\leq m\leq d$. Since $W$ is compact and $w=g(u)$ is proper, the number of all local coordinates can be taken to be finite. The property and application of RLCT is discussed in subsection 3.2.
@@
[…]
@@
holds without explicit values of $\lambda$ and $\nu(\beta)$, which shows that the difference between the generalization error and the training error is asymptotically equal to the posterior variance of point-wise log likelihood function. Hence we can compare their theoretical values with the experimental ones without their explicit values.
@@ 3.2 Real log canonical threshold
which was first created by Gel’fand [40] and has been studied by many mathematicians. By the definition, $\zeta(z)$ is an holomorphic function in $\Re(z)>0$ which can be analytically continued to the unique meromorphic function on the entire complex plane. The uniqueness of the analytic continuation can be proven by either the Hironaka resolution theorem in subsection 3.1 or the existence of Bernstein-Sato polynomial [41, 42]. There exist both a differential operator $D_{w}$ and a polynomial $b(z)$ such that, for arbitrary $z\in\mathbb{C}$ and $w\in W$,
@@
Then the monic polynomial that has the lowest order and satisfies this equation is called Bernstein-Sato polynomial. By Hironaka resolution theorem, All poles of the zeta function is real and negative numbers. The largest pole and its order of the zeta function are equal to $(-\lambda)$ and $m$ respectively. By using the zeta function, it is trivial that statistical learning theory of a pair
@@
    If there exists the unique $w_{0}\in W_{0}$ such that $\nabla^{2}L(w_{0})$ is positive definite and $\varphi(w_{0})>0$, then $\lambda=d/2$.
@@
By using these properties, RLCTs of important statistical models and learning machines were found by developing resolution procedures in neural networks [44, 45], deep linear networks [46], deep convolution ReLU network [47] with and without skip connection, normal mixtures [48], Poisson mixtures [49], multinomial mixtures [50], general and nonnegative matrix factorization [51, 52], Boltzmann machines [53], hidden and general Markov models [54, 55], and latent Dirichlet allocations [56].
In statistical inference, the free energy $F(1,X^{n})$ $(\beta=1)$ is often employed as an evaluation criterion of a pair $(p(x|w),\varphi(w))$. Because numerical calculation of the free energy requires heavy computational costs, several approximation methods were developed. RLCT is useful for such a purpose.
@@
where $\hat{w}$ is the parameter which minimizes $H(w,X^{n})$. In regular cases, the difference between the free energy and BIC is a constant order random variable. For general cases when $q(x)$ may be singular for $p(x|w)$, the singular BIC was proposed [58] by using the estimated RLCT $\hat{\lambda}$,
@@ 3.3 Singular Fluctuation
In this subsection, mathematical properties and applications of singular fluctuation are studied. The essential concept underlying singular fluctuation has been studied in many statistical contexts [61, 62, 63, 64].
@@
If $q(x)$ is regular for $p(x|w)$, the limit $\beta\rightarrow\infty$ results in the maximum likelihood method in statistics, in which posterior distribution converges to the delta function on the maximum likelihood estimator. Then
@@
holds, which was first found by Akaike [61]. The $2n$ times of the right hand side of this equation is called Akaike information criterion (AIC). If $q(x)$ is not realizable by $p(x|w)$, then $d$ is replaced by $\mbox{tr}(IJ^{-1})$ which is called by Takeuchi information criterion (TIC). They are defined as
${\rm AIC}$ $=T_{n}(\infty)+d/n,$ (23)
${\rm TIC}$ $=T_{n}(\infty)+\mbox{tr}(IJ^{-1})/n.$ (24)
Remark that, in practical applications, $2n$ times those values are used.
For the case when $\beta=1$ and $q(x)$ is regular for and realizable by a model, the $\mathbb{E}_{X^{n}}[G(1,X^{n})]$ can be estimated by deviance information criterion (DIC) [65]. The right hand side of eq.(19) is called WAIC which can be used to estimate the generalization error in general cases.
@@
Note that, if one has a numerical approximation software of the posterior average $\mathbb{E}_{w}^{\beta}[\;\;]$ by using Markov chain Monte Calro method, then $C_{n}(\beta)$ can be calculated by
@@
The difference between LOOCV and WAIC is smaller than any $1/n$ order random variable. In numerical calculation of eq.(25), the stability of the expectation $\mathbb{E}_{w}^{\beta}[p(X_{i}|w)^{-\beta}]$ is not ensured if a leverage sample point is contained in a sample [69, 70]. A new calculation method of LOOCV was proposed by using the approximation in MCMC distribution [71]. When the conditional probability $q(y|x)$ is estimated in regression problems, LOOCV requires that both $\{X_{i}\}$ and $\{(Y_{i}|X_{i})\}$ are independent, whereas information criteria do that only $\{(Y_{i}|X_{i})\}$ are independent. On the other hand, information criteria needs that the sample size $n$ is sufficiently large, hence LOOCV and information criteria are different statistical tools [36, 72]. The concept of generalization error was extended onto the problems such as covariate shift, causal inference [73], and overparametrized cases [74], and the generalized information criteria are now being developed.
@@ 3.4 Singular Learning Theory and AI Alignment
In 1966, the statistician I. J. Good, who firstly found the importance of the marginal likelihood $Z(1,X^{n})$ in statistical context, speculated as [75]
@@
Artificial neural network research started in the 1960s. Nowadays, artificial intelligence (AI) development is progressing so rapidly that it is difficult to predict to which direction AI takes us. It is said that alignment of artificial intelligence is necessary for safety and welfare of human being in the near future, where AI alignment is sometimes defined as AI working without deviating from the designer’s intention.
In order to consider AI alignment, we need all opinions from all different viewpoints, and problems to be studied are beging proposed [76, 77]. Note that singularities in learning machines make the free energy and the generalization loss smaller if Bayesian statistics is employed in machine learning, hence degenerate property of deep neural networks can be understood as an advantage for constructing artificial intelligence [27]. In AI alignment, we need to study developmental landscape of in-context learning in also transformers [78]. Studying deneracy in the loss function [79] is necessary in mechanistic interpretability [80]. The refined learning coefficient or RLCT is used for developmental interpretability [81].
In this subsection, we discuss AI alignment from the viewpoint of equivalence of statistical mechanics and machine learning. There are at least three problems caused by singularities in artificial neural networks: lack of identifiability, difficulty in design of the prior distribution, and phase transition in learning process. These three points originated from the hierarchical structure of learning machines. Remark that hierarchical structure is necessary for a learning machine to have an ability of efficient universal approximation, that is to say, an arbitrary continuous function can be approximated by multi-layered neural networks with higher precision. Moreover it makes the generalization error far smaller if a learning machine is over-parametrized. Therefore, solving the above three problems are necessary. I now explain these problems one by one.
The first problem is lack of identifiability. A statistical model or a leaning machine $p(x|w)$ is called identifiable if the map $w\mapsto p(x|w)$ is one-to-one. Many of classical statistical models are identifiable, hence the role of the parameter is uniquely determined. Therefore, by checking the obtained parameter, we can examine the reason why a machine answers an output for a given input. However, in singular learning machines, the role of the parameter is not uniquely determined, hence, even by checking the parameter, we cannot examine whether the model’s learning is subject to the designer’s intention.
@@
and preparing the quotient set $W\!/\!\sim$, the map from an equivalence class to a probability density function can be made one-to-one. This method is called blow-down in algebraic geometry. However, if $w_{0}$ is a singularity in the original parameter space, then its neighborhood in $W\!/\!\sim$ has infinite dimension in general, resulting that difficulty in understanding the parameter cannot be solved by considering $W\!/\!\sim$.
The second problem is difficulty in design of prior distribution. In the older Bayesianism in the 20th century, it was premised that uncertainty could be captured by a statistical model and the personal belief should be represented by a prior distribution. However, because of nonidentiability, each element of the parameter does not have any concrete meaning, hence the prior distribution cannot be determined by any a priori personal knowledge about data-generating process. One might think that Jeffreys’ prior may be chosen as an objective one, however, it makes the generalization error far larger than others [82], which destroys the advantage of hierarchical structure.
We need modern Bayesian statistics, in which a statistical model or a learning machine is only a candidate and that a prior distribution is not any belief of a person but a part of a statistical model. In the practical applications, ridge or lasso type prior distribution with hyperparameter is sometimes employed and the hyperparameter is optimized so that the generalization error or the free energy is made smaller by several criteria including cross validation.
The third problem is the phase transition phenomenon in learning process [83]. In classical statistical models, the optimal parameter $w_{0}$ is unique and the posterior distribution can be approximated as a normal distribution which converges to the delta function $\delta(w-w_{0})$ as the sample size increases and the posterior distribution may be understood as a gradual increase of confidence.
On the other hand, in deep learning, even if the sample size is huge, it is far smaller than the infinity for the deep neural network. Let $q(x)$ be unrealizable by $p(x|w)$. By using division of unity, there is a finite set of nonngegative $C_{0}^{\infty}$ class functions $\{\varphi_{k}(w)\}$ such that
@@
$F_{nk}(\beta)=nL(w_{k})+\frac{\lambda_{k}}{\beta}\log n+O(\log\log n).$
Then by using an inequality
$\min_{k}F_{nk}(\beta)-\log K\leq-\log\left(\sum_{k=1}^{K}\exp(-F_{nk}(\beta))\right)\leq\min_{k}F_{nk}(\beta),$
it follows that
@@
Therefore the free energy is given by the minimum of the local free energy. In statistical learning theory, $L(w_{k})$ and $\lambda_{k}\log n$ are called bias and variance, respectively. Note that $\exp(-F_{nk})$ is in proportion to the posterior probability of the local coordinate. Therefore the optimal local parameter set for minimum free energy is automatically chosen by the posterior probability. However, it is not equal to the optimal parameter set for the minimum generalization error. The local parameter set chosen by the posterior distribution jumps from a neighborhood to another one as sample size increases. Such jumping process makes it difficult to diagnose whether the learning process is out of order or not. This is a phase transition phenomenon with respect to the increase of the sample size.
Other phase transition phenomena can be also observed in different situations. In mixture models, a Dirichlet distribution is often chosen for a prior distribution of the mixture ratio. Then the posterior distribution has a phase transition according to the hyperparameter of Dirichlet distribution [36, 50]. In the different phases, RCLTs are different and the supports of asymptotic posterior distributions are different, which means that the local parameter sets for the different hyperparameters may be distant from each other [36].